Suppose that the following equations describe a patent-protected monopolist.

AR:      P = 4-0.0025(Q)

MR:     P = 4-0.005(Q)        

MC:     P = 0.4+0.003(Q)

If the firm is maximizing profits, what will their total revenue be?
Answer #1

Answer

The firm produces at MR=MC

equating both equations

4-0.005Q=0.4+0.003Q

0.008Q=3.6

Q=3.6/.008

Q=450

P=4-0.0025*450 ..... from AR curve

P=2.875

Total revenue =AR*Q

=2.875*450

=1293.75

The total revenue is $1293.75